### v3

# From Hypothesis to Publication: A Comprehensive Survey of AI-Driven Research Support Systems

Abstract
Research is a fundamental process driving the advancement of human civilization, yet it demands substantial time and effort from researchers. In recent years, the rapid development of artificial intelligence (AI) technologies has inspired researchers to explore how AI can accelerate and enhance research. To monitor relevant advancements, this paper presents a systematic review of the progress in this domain. Specifically, we organize the relevant studies into three main categories: hypothesis formulation, hypothesis validation, and manuscript publication. Hypothesis formulation involves knowledge synthesis and hypothesis generation. Hypothesis validation includes the verification of scientific claims, theorem proving, and experiment validation. Manuscript publication encompasses manuscript writing and the peer review process. Furthermore, we identify and discuss the current challenges faced in these areas, as well as potential future directions for research. Finally, we also offer a comprehensive overview of existing benchmarks and tools across various domains that support the integration of AI into the research process. We hope this paper serves as an introduction for beginners and fosters future research. Resources have been made publicly available [^1] https://github.com/zkzhou126/AI-for-Research.
From Hypothesis to Publication: A Comprehensive Survey of AI-Driven Research Support Systems
Zekun Zhou1, Xiaocheng Feng1,2thanks: Corresponding Author, Lei Huang1, Xiachong Feng3, Ziyun Song1, Ruihan Chen1, Liang Zhao1, Weitao Ma1, Yuxuan Gu1, Baoxin Wang4, Dayong Wu4, Guoping Hu4, Ting Liu1, Bing Qin1,2 1 Harbin Institute of Technology, Harbin, China  2 Peng Cheng Laboratory, Shenzhen, China 3 The University of Hong Kong, China  4 iFLYTEK Research, China {zkzhou,xcfeng,lhuang,zysong,rhchen,lzhao,wtma,yxgu,tliu,qinb}@ir.hit.edu.cn fengxc@hku.hk{bxwang2,dywu2,gphu}@iflytek.com
1 Introduction
Research is creative and systematic work aimed at expanding knowledge and driving civilization’s development (Eurostat, 2018). Researchers typically identify a topic, review relevant literature, synthesize existing knowledge, and formulate hypothesis, which are validated through theoretical and experimental methods. Findings are then documented in manuscripts that undergo peer review before publication (Benos et al., 2007; Boyko et al., 2023). However, this process is resource-intensive, requiring specialized expertise and posing entry barriers for researchers (Blaxter et al., 2010).
In recent years, artificial intelligence (AI) technologies, represented by large language models (LLMs), have experienced rapid development  (Brown et al., 2020; OpenAI, 2023; Dubey et al., 2024; Yang et al., 2024a; DeepSeek-AI et al., 2024; Guo et al., 2025). These models exhibit exceptional capabilities in text understanding, reasoning, and generation (Schaeffer et al., 2023). In this context, AI is increasingly involving the entire research pipeline (Messeri and Crockett, 2024), sparking extensive discussion about its implications for research  (Hutson, 2022; Williams et al., 2023; Morris, 2023; Fecher et al., 2023). Moreover, following the release of ChatGPT, approximately 20% of academic papers and peer-reviewed texts in certain fields have been modified by LLMs (Liang et al., 2024a, b). A study also reveals that 81% of researchers integrate LLMs into their workflows (Liao et al., 2024).
Figure 1: Overview of AI for research. The framework consists of three stages: hypothesis formulation, hypothesis validation, and manuscript publication. In the hypothesis formulation stage, knowledge integration leads to the proposal of an initial hypothesis after a topic is identified. The hypothesis validation stage involves verifying the hypothesis from three perspectives to ensure its correctness and validity. Finally, the manuscript publication stage focuses on drafting and publishing the validated hypothesis.
As the application of AI in research attracts increasing attention, a significant body of related studies has begun to emerge. To systematically synthesize existing research, we present comprehensive survey that emulates human researchers by using the research process as an organizing framework. Specifically, as depicted in Figure 1, the research process is divided into three key stages: (1) Hypothesis Formulation, involving knowledge synthesis and hypothesis generation; (2) Hypothesis Validation, encompassing scientific claim verification, theorem proving, and experiment validation; (3) Manuscript Publication, which focuses on academic publications and is further divided into manuscript writing and peer review.
Comparing with Existing Surveys Although Luo et al. (2025) reviews the application of AI in research, it predominantly focuses on LLMs, while neglecting the knowledge synthesis that precedes hypothesis generation and the theoretical validation of hypothesis. Other surveys concentrate on more specific areas, such as paper recommendation (Beel et al., 2016; Bai et al., 2019; Kreutz and Schenkel, 2022), scientific literature review (Altmami and Menai, 2022), hypothesis generation (Kulkarni et al., 2025), scientific claim verification (Vladika and Matthes, 2023; Dmonte et al., 2024), theorem proving (Li et al., 2024e), manuscript writing (Li and Ouyang, 2024), and peer review  (Lin et al., 2023a; Kousha and Thelwall, 2024). Additionally, certain surveys emphasize the application of AI in scientific domains (Zheng et al., 2023b; Zhang et al., 2024d; Gridach et al., 2025).
Contributions Our contributions can be summarized as follows: (1) We align the relevant fields with the research process of human researchers, systematically integrating and extending these aspects while primarily focusing on the research process itself. (2) We introduce a meticulous taxonomy (shown in Figure 2). (3) We provide a summary of tools that can assist in the research process. (4) We discuss new frontiers, outline their challenges, and shed light on future research.
{forest}
forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8em,font=,, where level=2text width=13em,font=,, where level=3text width=11.5em,font=,, where level=4text width=13em,font=,, [ A survey of AI for Research, ver [ Hypothesis
Formulation  (§2) [ Knowledge Synthesis  (§2.1) [ Research Paper
Recommendation  (§2.1.1) [ E.g.,  ComLittee Kang et al. (2023) , PaperWeaver Lee et al. (2024) ,leaf, text width=34.5em ] ] [ Systematic
Literature Review  (§2.1.2) [ E.g.,  AutoSurvey  Wang et al. (2024e) , OpenScholar  Asai et al. (2024) ,leaf, text width=34.5em ] ] ] [ Hypotheses Generation  (§2.2) [ E.g.,  SciMON  Wang et al. (2024c) , MOOSE  Yang et al. (2024b) , Dolphin  Yuan et al. (2025) ,leaf, text width=47.6em ] ] ] [ Hypothesis
Validation  (§3) [ Scientific
Claim Verification  (§3.1) [ E.g.,  FactKG  Kim et al. (2023) , SFAVEL Bazaga et al. (2024) , ClaimVer Dammu et al. (2024) ,leaf, text width=47.6em ] ] [ Theorem Proving  (§3.2) [ E.g.,  Dt-solver  Wang et al. (2023b) , LEGO-Prover  Wang et al. (2024b) , DeepSeek-Prover  Xin et al. (2024) ,leaf, text width=47.6em ] ] [ Experiment Validation  (§3.3) [ E.g.,  SANDBOX  Liu et al. (2023b) , CRISPR-GPT  Huang et al. (2024a) , MLR-Copilot  Li et al. (2024d) ,leaf, text width=47.6em ] ] ] [ Manuscript
Publication  (§4) [ Manuscript Writing  (§4.1) [ E.g.,  Scilit  Gu and Hahnloser (2023) , UR3WG  Shi et al. (2023) , STEP-BACK  Tang et al. (2024a) ,leaf, text width=47.6em ] ] [ Peer Review  (§4.2) [ E.g.,  SWIF2T  Chamoun et al. (2024) , GLIMPSE  Darrin et al. (2024) , MetaWriter  Sun et al. (2024) ,leaf, text width=47.6em ] ] ] ]
Figure 2: Taxonomy of Hypothesis Formulation, Hypothesis Validation and Manuscript Publication (This is a simplified version, full version in Figure 6).
Survey Organization We first elaborate hypothesis formulation (§2), followed by hypothesis validation (§3) and manuscript publication (§4). Additionally, we present benchmarks (§5), and tools (§6) that utilized in research. Finally, we outline challenges as well as future directions (§7). In the Appendix, we provide further discussion on open questions (§A), challenges faced in different domains (§B), discusssion about relevant ethical considerations (§C), and a comparison of capabilities among different methods (§D).
2 Hypothesis Formulation
This stage centers on the process of hypothesis formulation. As illustrated in Figure 3, it commences with developing a comprehensive understanding of the domain, followed by identifying a specific aspect and generating pertinent hypothesis. This section is further structured into two key components: Knowledge Synthesis and Hypothesis Generation.
2.1 Knowledge Synthesis
Knowledge synthesis constitutes the foundational step in the research process. During this phase, researchers are required to identify and critically evaluate existing literature to establish a thorough understanding of the field. This step is pivotal for uncovering new research directions, refining methodologies, and supporting evidence-based decision-making (Asai et al., 2024). In this section, the process of knowledge synthesis is divided into two modules: Research Paper Recommendation and Systematic Literature Review.
2.1.1 Research Paper Recommendation
Research paper recommendation (RPR) identifies and recommends novel and seminal articles aligned with researchers’ interests. Prior studies have shown that recommendation systems outperform keyword-based search engines in terms of efficiency and reliability when extracting valuable insights from large-scale datasets (Bai et al., 2019). Existing methodologies are broadly categorized into four paradigms: content-based filtering, collaborative filtering, graph-based approaches, and hybrid systems (Beel et al., 2016; Li and Zou, 2019; Bai et al., 2019; Shahid et al., 2020). Recent advancements propose multi-dimensional classification frameworks based on data source utilization  (Kreutz and Schenkel, 2022).
Recent trends in research suggest a decline in publication volumes related to RPR  (Sharma et al., 2023), alongside an increasing focus on user-centric optimizations. Existing studies emphasize the limitations of traditional paper-centric interaction models and advocate for more effective utilization of author relationship graphs  (Kang et al., 2023). Multi-stage recommendation architectures that integrate diverse methodologies have been shown to achieve superior performance  (Pinedo et al., 2024; Stergiopoulos et al., 2024). Visualization techniques that link recommended papers to users’ publication histories via knowledge graphs  (Kang et al., 2022) and LLMs-powered comparative analysis frameworks  (Lee et al., 2024) represent emerging directions for enhancing interpretability and contextual relevance.
2.1.2 Systematic Literature Review
Systematic literature review (SLR) constitutes a rigorous and structured methodology for evaluating and integrating prior research on a specific topic  (Webster and Watson, 2002; Zhu et al., 2023; Bolaños et al., 2024). In contrast to single-document summaries (Elhadad et al., 2005), SLR entails synthesizing information across multiple related scientific documents (Altmami and Menai, 2022). SLR can further be divided into two stages: outline generation and full-text generation (Shao et al., 2024; Agarwal et al., 2024b; Block and Kuckertz, 2024).
Outline generation, especially structured outline generation, is highlighted by recent studies as a pivotal factor in enhancing the quality of surveys.  Zhu et al. (2023) demonstrated that hierarchical frameworks substantially enhance survey coherence. AutoSurvey  (Wang et al., 2024e) extended conventional outline generation by recommending both sub-chapter titles and detailed content descriptions, ensuring comprehensive topic coverage. Additionally, multi-level topic generation via clustering methodologies has been proposed as an effective strategy for organizing survey structures  (Katz et al., 2024). Advanced systems such as STORM  (Shao et al., 2024) employed LLM-driven outline drafting combined with multi-agent discussion cycles to iteratively refine the generated outlines. Tree-based hierarchical architectures have gained increasing adoption in this domain. For instance, CHIME  (Hsu et al., 2024) optimized LLM-generated hierarchies through human-AI collaboration, while HiReview  (Hu et al., 2024b) demonstrated the efficacy of multi-layer tree representations for systematic knowledge organization.
Full-text generation follows the outline generation stage. AutoSurvey and  Lai et al. (2024) utilized LLMs with carefully designed prompts to construct comprehensive literature reviews step-by-step. PaperQA2  (Skarlinski et al., 2024) introduced an iterative agent-based approach for generating reviews, while STORM employed multi-agent conversation data for this purpose. LitLLM  (Agarwal et al., 2024a) and  Agarwal et al. (2024b) adopted a plan-based search enhancement strategy. KGSum  (Wang et al., 2022a) integrated knowledge graph information into paper encoding and used a two-stage decoder for summary generation. Bio-SIEVE  (Robinson et al., 2023) and  Susnjak et al. (2024) fine-tuned LLMs for automatic review generation. OpenScholar  (Asai et al., 2024) developed a pipeline that trained a new model without relying on a dedicated survey-generation model.
Figure 3: This figure illustrates the hypothesis formulation process, consisting of two stages: knowledge synthesis and hypothesis generation, which together produce an initial hypothesis related to a specific topic.
2.2 Hypothesis Generation
Hypothesis generation, known as idea generation, refers to the process of coming up with new concepts, solutions, or approaches. It is the most important step in driving the progress of the entire research  (Qi et al., 2023).
Early work focused more on predicting relationships between concepts, because researchers believed that new concepts come from links with old concepts (Henry and McInnes, 2017; Krenn et al., 2022). As language models became more powerful (Zhao et al., 2023a), researchers are beginning to focus on open-ended idea generation  (Girotra et al., 2023; Si et al., 2024; Kumar et al., 2024). Recent advancements in AI-driven hypothesis generation highlight diverse approaches to research conceptualization. For instance, MOOSE-Chem  (Yang et al., 2024c) and IdeaSynth  (Pu et al., 2024) used LLMs to bridge inspiration-to-hypothesis transformation via interactive frameworks. The remaining research primarily falls into two areas: enhancing input data quality and improving the quality of generated hypothesis.
Input data quality improvement is demonstrated by Majumder et al. (2024a); Liu et al. (2024a), who showed that LLMs can generate comprehensive hypothesis from existing academic data. Literature organization strategies have evolved through various methodologies, including triplet representations  (Wang et al., 2024c), chain-based architectures  (Li et al., 2024a), and complex database systems  (Wang et al., 2024d). Knowledge graphs emerge as critical infrastructure (Hogan et al., 2021), enabling semantic relationship mapping via subgraph identification (Buehler, 2024; Ghafarollahi and Buehler, 2024). Notably, SciMuse  (Gu and Krenn, 2024) pioneered researcher-specific hypothesis generation by constructing personalized knowledge graphs.
Hypothesis quality improvement has been addressed through feedback and iteration (Rabby et al., 2025), as proposed by HypoGeniC  (Zhou et al., 2024) and MOOSE  (Yang et al., 2024b). Feedback mechanisms include direct responses to hypothesis  (Baek et al., 2024), experimental outcome evaluations  (Ma et al., 2024; Yuan et al., 2025), comparison with the existing literature (Schmidgall and Moor, 2025), and automated peer review comments  (Lu et al., 2024). FunSearch (Romera-Paredes et al., 2024) generates solutions by iteratively combining the innovative capabilities of LLM with the verification capabilities of an evaluator. Beyond iterative feedback, collaborative efforts among researchers have also been recognized, leading to multi-agent hypothesis generation approaches  (Nigam et al., 2024; Ghafarollahi and Buehler, 2024). VIRSCI  (Su et al., 2024) further optimized this process by customizing knowledge for each agent. Additionally, the Nova framework  (Hu et al., 2024a) was introduced to refine hypothesis by leveraging outputs from other research as input.
Knowledge synthesis and hypothesis generation comprise the hypothesis formulation phase. Research paper recommendation supports knowledge acquisition, while systematic literature review aids organization within knowledge synthesis. Recent advances integrate LLMs  (de la Torre-López et al., 2023) to enhance knowledge integration  (Huang and Tan, 2023; Gupta et al., 2023; Kacena et al., 2024; Tang et al., 2024b). By developing a deep understanding of a domain through knowledge synthesis, researchers can identify research directions and use hypothesis generation techniques to formulate hypothesis. Additionally, the distinction between scientific discovery and hypothesis generation is discussed in  §A.
3 Hypothesis Validation
In scientific research, any proposed hypothesis must undergo rigorous validation to establish its validity. In some studies, this process is also referred to as ’falsification’ (Liu et al., 2024d; Huang et al., 2025). As illustrated in Figure  4, this section explores the application of AI in verifying scientific hypothesis through three methodological components: Scientific Claim Verification, Theorem Proving, and Experiment Validation.
Figure 4: This figure illustrates the various perspectives for hypothesis validation during the hypothesis validation stage. A hypothesis is typically divided into scientific claims and theorems, with SCI-claim verification (scientific claim verification) and theorem proving ensuring theoretical correctness, while experiment validation assesses practical feasibility.
3.1 Scientific Claim Verification
Scientific claim verification, also referred to as scientific fact-checking or scientific contradiction detection, aims to assess the veracity of claims related to scientific knowledge. This process assists scientists in verifying research hypothesis, discovering evidence, and advancing scientific work  (Wadden et al., 2020; Vladika and Matthes, 2023; Skarlinski et al., 2024). Research on scientific claim verification primarily focuses on three key elements: the claim, the evidence, and the validity of the claim  (Dmonte et al., 2024).
Claim
Studies have highlighted that certain claims lack supporting evidence  (Wührl et al., 2024a), while others have demonstrated the ability to perform claim-evidence alignment without annotated data  (Bazaga et al., 2024). Additionally, methods such as HiSS  (Zhang and Gao, 2023) and ProToCo  (Zeng and Gao, 2023) proposed generating multiple claim variants to enhance verification.
Evidence
Existing research has explored various aspects related to evidence, including evidentiary sources  (Vladika and Matthes, 2024a), retrieval configurations  (Vladika and Matthes, 2024b), strategies for identifying and mitigating flawed evidence  (Glockner et al., 2022; Wührl et al., 2024b; Glockner et al., 2024a), and approaches for processing sentence-level  (Pan et al., 2023b) versus document-level indicators  (Wadden et al., 2022b).
Verification
In the verification results generation phase, MAGIC (Kao and Yen, 2024) and SERIf (Cao et al., 2024b) proposed utilizing LLMs to synthesize evidence into more comprehensive information. FactKG  (Kim et al., 2023) and  Muharram and Purwarianti (2024) structured evidence into knowledge graphs, enabling claim attribution  (Dammu et al., 2024; Wu et al., 2023). Furthermore,  Atanasova et al. (2020); Krishna et al. (2022); Pan et al. (2023a); Eldifrawi et al. (2024); Zhang et al. (2024b) advocated for generating explanatory annotations alongside experimental outcomes during the verification process. Meanwhile,  Das et al. (2023); Altuncu et al. (2023) emphasized the critical role of domain expertise in ensuring accurate verification.
3.2 Theorem Proving
Theorem proving constitutes a subtask of logical reasoning, aimed at reinforcing the validity of the underlying theory within a hypothesis  (Pease et al., 2019; Yang et al., 2023c; Li et al., 2024e).
Following the proposal of GPT-f  (Polu and Sutskever, 2020) to utilize generative language models for theorem proving, researchers initially combined search algorithms with language models  (Lample et al., 2022; Wang et al., 2023b). However, a limitation in search-based approaches was later identified by  Wang et al. (2024a), who highlighted their tendency to explore insignificant intermediate conjectures. This led some teams to abandon search algorithms entirely. Subsequently, alternative methods emerged, such as the two-stage framework proposed by Jiang et al. (2023) and Lin et al. (2024), which prioritized informal conceptual generation before formal proof construction. Thor  (Jiang et al., 2022a) introduced theorem libraries to accelerate proof generation, an approach enhanced by Logo-power  (Wang et al., 2024b) through dynamic libraries. Studies like Baldur  (First et al., 2023), Mustard  (Huang et al., 2024c), and DeepSeek-Prover  (Xin et al., 2024) demonstrated improvements via targeted data synthesis and fine-tuning, though COPRA  (Thakur et al., 2024) questioned their generalizability and proposed an environment-agnostic alternative. Complementary strategies included theoretical decomposition into sub-goals  (Zhao et al., 2023b) and leveraging LLMs as collaborative assistants in interactive environments  (Song et al., 2024).
3.3 Experiment Validation
Experiment validation involves designing and conducting experiments based on the hypothesis. The empirical validity of the hypothesis is then determined through analysis of the experimental results (Huang et al., 2024b).
Experiment validation represents a time-consuming component of scientific research. Recent advancements in LLMs have enhanced their ability to plan and reason about experimental tasks  (Kambhampati et al., 2024), prompting researchers to use these models for designing and implementing experiments  (Ruan et al., 2024b). To ensure accuracy, studies such as  Zhang et al. (2023) and  Arlt et al. (2024) imposed input/output constraints, though this reduced generalizability. To address this,  Boiko et al. (2023); Bran et al. (2024); Huang et al. (2024a) integrated tools to expand model capabilities. Full automation was achieved by  Ni and Buehler (2023); Li et al. (2024a); Lu et al. (2024) through prompt-guided multi-agent collaboration.  Madaan et al. (2023); Yuan et al. (2025) further highlighted that the integration of feedback mechanisms demonstrated potential for enhancing design quality, while  Zhang et al. (2024a); Liu et al. (2024c); Ni et al. (2024) employed experimental outcomes to refine hyperparameter configurations, and  Szymanski et al. (2023); Li et al. (2024d); Baek et al. (2024) leveraged agent-generated analytical insights to facilitate iterative hypothesis refinement. In contrast, social science research often uses LLMs as experimental subjects to simulate human participants  (Liu et al., 2023b; Manning et al., 2024; Mou et al., 2024).
A hypothesis can be conceptualized as consisting of two key components: claims and theorems. Scientific claim verification and theorem proving offer theoretical validation of hypothesis through formal reasoning and logical deduction, whereas experiment validation provides comprehensive practical validation via empirical testing.
4 Manuscript Publication
Upon validating a hypothesis as feasible, researchers generally progress to the publication stage. As depicted in Figure 5, this section categorizes Manuscript Publication into two primary components: Manuscript Writing and Peer Review.
4.1 Manuscript Writing
Manuscript writing, also referred to as scientific or research writing. At this stage, researchers articulate the hypothesis they have formulated and the results they have validated in the form of a scholarly paper. This process is crucial, as it not only disseminates findings but also deepens researchers’ understanding of their work  (Colyar, 2009).
Early shared tasks focused on assisting researchers in writing or analyzing linguistic features (Dale and Kilgarriff, 2010; Daudaravicius, 2015). Recent advances have led to three main directions: citation text generation, related work generation, and complete manuscript generation.
Figure 5: This figure shows the transformation of a validated hypothesis into a publication, leveraging outputs from the hypothesis formulation and validation stages.
Citation Text Generation (Sentence Level)
A subset of research on AI in scientific writing has focused on citation text generation, which addresses the academic need for referencing prior work while mitigating model inaccuracies  (Gao et al., 2023b; Gu and Hahnloser, 2023). For instance,  Wang et al. (2022b) developed an automated citation generation system by integrating manuscript content with citation graphs. However, its reliance on rigid template-based architectures led to inflexible citation formats. This limitation motivated subsequent studies to propose incorporating citation intent as a control parameter during text generation, aiming to improve contextual relevance and rhetorical adaptability  (Yu et al., 2022; Jung et al., 2022; Koo et al., 2023; Gu and Hahnloser, 2024).
Related Work Generation (Paragraph Level)
In contrast to citation text generation, several studies have focused on related work generation in scholarly writing, emphasizing the production of multiple citation texts and the systematic analysis of inter-citation relationships  (Li and Ouyang, 2022, 2024). The ScholaCite framework  (Martin-Boyle et al., 2024) leveraged GPT-4 to cluster citation sources and generate draft literature review sections, although it required manually provided reference lists. By contrast, the UR3WG system  (Shi et al., 2023) adopted a retrieval-augmented architecture integrated with large language models to autonomously acquire relevant references. To improve the quality of generated related work sections,  Yu et al. (2024b) utilized GNNs to model complex relational dynamics between target manuscripts and cited literature, while  Nishimura et al. (2024) initiative advocated for explicit novelty assertions regarding referenced publications.
Complete Manuscripts Generation (Full-text Level)
The aforementioned investigations primarily focused on specific components of scientific writing, while a study by  Lai et al. (2024) explored the progressive generation of complete manuscripts via structured workflows. The AI-Scientist system  (Lu et al., 2024) further introduced section-wise self-reflection mechanisms to enhance compositional coherence. Several studies emphasized human-AI collaborative frameworks for improving writing efficiency  (Lin, 2024; Feng et al., 2024; Ifargan et al., 2024), whereas  Tang et al. (2024a) concentrated on enabling personalized content generation in multi-author collaborative environments. Following initial manuscript drafting, subsequent text revision processes were systematically examined  (Du et al., 2022b; Jourdan et al., 2023; Dang et al., 2025). The OREO system  (Li et al., 2022) utilized attribute classification for iterative in-situ editing, while  Du et al. (2022a); Pividori and Greene (2024) incorporated researcher feedback loops for progressive text optimization. Notably,  Kim et al. (2022); Chamoun et al. (2024); D’Arcy et al. (2024b) proposed replacing manual feedback with automated evaluation metrics.
4.2 Peer Review
Peer review serves as a critical mechanism for improving the quality of academic manuscripts through feedback and evaluation, forming the cornerstone of quality control in scientific research. However, the process is hindered by its slow pace, high time consumption, and increasing strain due to the growing academic workload  (Lin et al., 2023a; Kousha and Thelwall, 2024; Thelwall and Yaghi, 2024). To address these challenges and enhance manuscript quality, researchers have investigated the application of AI in peer review  (Yuan et al., 2022; Liu and Shah, 2023; Niu et al., 2023; Kuznetsov et al., 2024; Thakkar et al., 2025). Peer review can be categorized into two main types: paper review generation and meta-review generation.
Paper Review Generation
In paper review generation, reviewers provide both scores and evaluations for manuscripts. For instance,  Setio and Tsuchiya (2022) formulated score prediction as a regression task,  Muangkammuen et al. (2022) utilized semi-supervised learning, and  Couto et al. (2024) treated the task as a classification problem to evaluate the alignment between manuscripts and review criteria. While these approaches focused on label prediction for paper reviews,  Yuan and Liu (2022) extended the scope by directly generating reviews through the construction of a concept graph integrated with a citation graph.
Subsequently, a pilot study conducted by  Robertson (2023) demonstrated the capability of GPT-4 to generate paper reviews. Further investigations, such as those by AI-Scientist  (Lu et al., 2024) and  Liang et al. (2023), evaluated its performance as a review agent. Additionally, systems like MARG  (D’Arcy et al., 2024a) and SWIF2T  (Chamoun et al., 2024) employed multi-agent frameworks to generate reviews via internal discussions and task decomposition. In contrast, AgentReview  (Jin et al., 2024) and  Tan et al. (2024) modeled the review process as a dynamic, multi-turn dialogue. Furthermore, CycleResearcher  (Weng et al., 2024) and OpenReviewer  (Idahl and Ahmadi, 2024) fine-tuned models for comparative reviews and structured outputs aligned with conference guidelines.
Meta-Review Generation
In meta-review generation, chairs are tasked with identifying a paper’s core contributions, strengths, and weaknesses while synthesizing expert opinions on manuscript quality. Meta-reviews are conceptualized as abstractions of comments, discussions, and paper abstracts  (Li et al., 2023).  Santu et al. (2024) investigated the use of LLMs for automated meta-review generation, while  Zeng et al. (2023) proposed a guided, iterative prompting approach. MetaWriter  (Sun et al., 2024) utilized LLMs to extract key reviewer arguments, whereas GLIMPSE  (Darrin et al., 2024) and  Kumar et al. (2023) focused on reconciling conflicting statements to ensure fairness. Additionally,  Li et al. (2024b) introduced a three-layer sentiment consolidation framework for meta-review generation, and PeerArg  (Sukpanichnant et al., 2024) integrated LLMs with knowledge representation to address subjectivity and bias via a multiparty argumentation framework (MPAF). DeepReview (Zhu et al., 2025) generates a comprehensive meta-review by simulating expert evaluation across multiple dimensions.
During the Manuscript Publication phase, researchers can leverage AI to systematically complete manuscript writing by incorporating validated hypothesis, related papers, and literature reviews. The manuscript is subsequently subjected to peer review, involving iterative revisions before culminating in its final publication.
5 Benchmarks
Given that AI for research spans multiple disciplines, the tasks addressed within each domain vary significantly. To facilitate cross-domain exploration, we provide a summary of benchmarks associated with various areas, including research paper recommendation, systematic literature review, hypothesis generation, scientific claim verification, theorem proving, experiment verification, manuscript writing, and peer review. An overview of these benchmarks is presented in Table 9.
6 Tools
To accelerate the research workflow, we have curated a collection of tools designed to support various stages of the research process, with their applicability specified for each stage. To ensure practical relevance, our selection criteria emphasize tools that are publicly accessible or demonstrate significant influence on GitHub. A comprehensive overview of these tools is presented in Table 10.
7 Challenges
We identify several intriguing and promising avenues for future research.
7.1 Integration of Diverse Research Tasks
The research process is an integrated pipeline of interdependent stages. Paper recommendation and literature review provide an AI tool with a field overview and relevant works, ensuring that hypothesis generation is informed and of higher quality. Hypothesis validation assesses feasibility both logically and practically, with results feeding back to refine the hypothesis. In manuscript writing, validated hypotheses and prior outputs serve as key inputs. Peer review evaluates the manuscript and offers feedback across modules, enabling the hypothesis generator to adjust content accordingly (Lu et al., 2024). In addition, combinations can also be made between some small fields, for instance, meta-review generation could be integrated with scientific claim verification, experiment verification could be linked with hypothesis formulation (Jansen et al., 2025; Yuan et al., 2025; Liu et al., 2024d), and research paper recommendation systems could be connected with manuscript writing processes (Gu and Hahnloser, 2023). Furthermore, some studies have begun to emphasize the development of systems capable of covering multiple stages of the research process (Jansen et al., 2024; Weng et al., 2024; Yu et al., 2024a).
7.2 Integration with Reasoning-Oriented Language Models
Research is a process that places a significant emphasis on logic and reasoning. Theorem proving serves as a subtask within logical reasoning (Li et al., 2024e), while hypothesis generation is widely recognized as the primary form of reasoning employed by scientists when observing the world and proposing hypothesis to explain these observations (Yang et al., 2024b). Experiment verification, in turn, demands a high degree of planning capability from models (Kambhampati et al., 2024). Recent advances in reasoning-oriented language models, such as OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025), have substantially enhanced the reasoning abilities of these models. Consequently, we posit that integrating reasoning language models with reasoning tasks is a promising future direction. This prediction was validated by experiments conducted by  Schmidgall et al. (2025) using o1-Preview.
Furthermore, in Appendix  §B, we provide a summary of the challenges in hypothesis formulation, validation, and manuscript publication.
8 Conclusion
This paper provides a systematic survey of existing research on AI for research, offering a comprehensive review of the advancements in the field. Within each category, we offer detailed descriptions of the associated subfields. In addition, we examine current challenges, ethical considerations, and potential directions for future research. To support researchers in exploring AI-driven research applications and enhancing workflow efficiency, we also summarize existing benchmarks and tools, accompanied by a comparative analysis of representative methods and their capabilities.
Furthermore, in the course of investigating various subfields within AI for research, we observed that this domain remains in its infancy. Research in numerous directions remains at an experimental stage, and substantial progress is necessary before these approaches can be effectively applied in practical scenarios. We hope that this survey serves as an introduction to the field for researchers and contributes to its continued advancement.
Limitation
This study presents a comprehensive survey of AI for research, based on the framework of the research process conducted by human researchers.
We have made our best effort, but there may still be some limitations. Due to space constraints, we provide only concise summaries of each method without detailed technical elaboration. Given the rapid progress in AI and the expanding research landscape, we primarily focus on works published after 2022, with earlier studies receiving less attention. To emphasize areas that closely mimic the human research process, some topics are excluded from the main text but briefly discussed in Appendix §A. Moreover, as AI for Research is still an emerging field, the lack of standardized benchmarks and evaluation metrics hinders direct comparison. Nonetheless, we offer a comparative analysis of representative methods across domains using attribute graphs in Appendix §D.
Appendix A Further Discussion
Open Question: What is the difference between AI for science and AI for research?
We posit that AI for research constitutes a subset of AI for science. While AI for research primarily focuses on supporting or automating the research process, it is not domain-specific and places greater emphasis on methodological advancements. In contrast, AI for science extends beyond the research process to include result-oriented discovery processes within specific domains, such as materials design, drug discovery, biology, and the solution of partial differential equations (Zheng et al., 2023b; AI4Science and Quantum, 2023; Zhang et al., 2024d).
Open Question: What is the difference between hypothesis generation and scientific discovery?
Hypothesis generation, which is primarily based on literature-based review (LBD) (Swanson, 1986; Sebastian et al., 2017), emphasizing the process by which researchers generate new concepts, solutions, or approaches through existing research and their own reasoning. Scientific discovery encompasses not only hypothesis generation, but also innovation in fields like molecular optimization and drug development  (Ye et al., 2024; Liu et al., 2024b), driven by outcome-oriented results.
Open Question: What is the difference between systematic literature review and related work generation?
Existing research frequently addresses the systematic literature survey, which constitutes a component of the knowledge synthesis process during hypothesis formulation, alongside the related work generation phase in manuscript writing (Luo et al., 2025). However, we argue that these two tasks are distinct in nature. The systematic literature survey primarily focuses on summarizing knowledge extracted from diverse scientific documents, thereby assisting researchers in acquiring an initial understanding of a specific field (Altmami and Menai, 2022). In contrast, related work generation focuses on the writing process, emphasizing selection of pertinent literature and effective content structuring  (Nishimura et al., 2024).
Discussion: Potential links between artificial intelligence systems and human research practice
  • In research paper recommendation, PaperWeaver (Lee et al., 2024) offers an interactive page that allows users to modify the topics they are interested in.
  • In systematic literature review,  Block and Kuckertz (2024) highlights the significant role of humans, including setting correct questions and individualized problem-solving and theorizing. Meanwhile,  Hsu et al. (2024) emphasizes manual correction during the outline generation process.
  • In hypothesis generation, AI engages more closely with human researchers, ranging from scenarios where humans provide the core ideas and AI contributes by iteratively refining them  (Pu et al., 2024), to more collaborative settings where humans and AI engage in dialogue to facilitate new scientific discoveries  (Ni et al., 2024; Liu et al., 2024b; Ye et al., 2024).
  • In scientific claim verification,  Altuncu et al. (2023); Das et al. (2023) highlight the critical role of experts in countering fake scientific news and advocate for the incorporation of expert opinions as a form of evidence.
  • In theorem proving,  Song et al. (2024) proposes leveraging LLMs as assistants to human researchers by generating suggested proof steps throughout the proving process.
  • In experiment Validation,  Ni et al. (2024) enhances the experimental setup through human-AI dialogue, whereas  Li et al. (2024d) incorporates human input and real-time adjustments during the execution phase to optimize experimental design.
  • In manuscript writing,  Ifargan et al. (2024); Feng et al. (2024); Du et al. (2022a) require human intervention to suggest improvements to AI-generated paragraphs and enhance their quality through interactive methods.
  • In peer review,  Kumar et al. (2023); Darrin et al. (2024) advocate for assigning the responsibility of generating meta-reviews to human researchers. The role of AI is to assist by identifying conflicts among reviewers’ opinions and supporting the chair in the scoring process, rather than independently assigning scores.
Discussion: The involvement of AI in manuscript writing
The application of AI in manuscript writing has been accompanied by significant controversy. As LLMs demonstrated advanced capabilities, an increasing number of researchers began adopting these systems for scholarly composition  (Liang et al., 2024b; Gao et al., 2023a). This trend raised concerns within the academic community  (Salvagno et al., 2023), with scholars explicitly opposing the attribution of authorship to AI systems  (Lee, 2023). Despite these reservations, the substantial time efficiencies offered by this technology led researchers to gradually accept AI-assisted writing practices  (Gruda, 2024; Huang and Tan, 2023; Chen, 2023).This shift ultimately led to formal guidelines issued by leading academic journals  (Ganjavi et al., 2024; Xu, 2025).
Discussion: Some areas that have not been discussed
In addition to the eight areas discussed above, there are other lines of work that also aim to support scientific research, such as reading assistance (Kang et al., 2020; Head et al., 2021; Lo et al., 2023), which helps researchers read academic papers; literature processing [^2] https://sdproc.org, which handles documents in various formats to provide effective data for subsequent tasks; as well as code and data generation (Bauer et al., 2024; Zheng et al., 2023a), which serve as a foundation for experimental validation. However, as our focus is on the core process of scientific research, we have chosen to omit these aspects from the main text.
Appendix B Challenges
B.1 Hypothesis Formulation
Knowledge Synthesize
Existing paper recommendation tools predominantly rely on the metadata of existing publications to suggest related articles, which often results in a lack of user-specific targeting and insufficiently detailed presentation that hampers comprehension. Leveraging LLMs can facilitate the construction of dynamic user profiles, enabling personalized literature recommendations and enhancing the richness of the contextual information provided for each recommended article, ultimately improving the user experience. In the process of generating systematic literature reviews, our practical experience reveals that the outline generation tools often produces redundant results with insufficient hierarchical structure. Moreover, the full-text generation process is prone to hallucinations—for instance, statements may not correspond to the cited articles—a pervasive issue in large language models (Huang et al., 2023; Bolaños et al., 2024; Susnjak et al., 2024). This problem can be ameliorated by enhancing the foundational model capabilities or by incorporating citation tracing.
Hypothesis Generation
Most existing tools generate hypotheses by designing prompts or constructing systematic frameworks, which heavily rely on the capabilities of pre-trained models. However, these methods struggle to balance the novelty, feasibility, and validity of the hypotheses (Li et al., 2024c).Furthermore, our investigation reveals that many current approaches adopt novelty and feasibility as evaluation metrics; these metrics are either difficult to quantify or require manual scoring, which can introduce bias. To date, there is no unified benchmark to compare the various methods, and we believe that future research should prioritize establishing a unified metric that objectively reflects the strengths and weaknesses of different approaches.
B.2 Hypothesis Validation
Most existing scientific claim verification tools are largely confined to specific domains, exhibiting poor generalizability, which limits their practical applicability  (Vladika and Matthes, 2023). Theorem proving, the scarcity of relevant data adversely affects performance improvements through training , results across different proof assistants are not directly comparable, and the lack of standardized evaluation benchmarks presents numerous challenges. Moreover, current approaches remain predominantly in the research stage and lack practical tools that facilitate interaction with researchers  (Li et al., 2024e). Experiment Validation, as automatically generated experiments often suffer from a lack of methodological rigor, practical feasibility, and alignment with the original research objectives  (Lou et al., 2024). All these fields require rigorous logical reasoning, and I believe that the recent surge in advanced reasoning technologies could potentially address these issues.
B.3 Manuscript Publication
Similar to systematic literature surveys, manuscript writing is also adversely affected by hallucination issues (Athaluri et al., 2023; Huang et al., 2023). Even when forced citation generation is employed, incorrect references may still be introduced (Aljamaan et al., 2024). Furthermore, the text generated by models requires meticulous examination by researchers to avoid ethical concerns, such as plagiarism risks (Salvagno et al., 2023). AI-generated manuscript reviews frequently provide vague suggestions and are susceptible to biases (Chamoun et al., 2024; Drori and Te’eni, 2024). Additionally, during meta-review generation, models can be misled by erroneous information arising from the manuscript review process (Kumar et al., 2023). To address these issues, it may be necessary for the industry to establish appropriate regulations or to employ AI-based methods for detecting AI-generated papers and reviews (Lin et al., 2023a).
Appendix C Ethical Considerations
AI has demonstrated significant potential in enhancing productivity by mitigating human limitations, thereby motivating increased investigation into its capacity to accelerate the research process  (Messeri and Crockett, 2024). Nevertheless, its integration into scientific workflows introduces a range of ethical concerns  (Fecher et al., 2023; Morris, 2023), including algorithmic biases, data privacy issues, risks of plagiarism, and the broader implications of AI-generated content for research communities. In this work, we examine these ethical challenges across the key stages of the research lifecycle: hypothesis formulation, validation, and publication.
During the hypothesis formulation stage, research paper recommendation systems and literature reviews are commonly employed; however, they often suffer from limitations that can lead to the formation of information bubbles and restrict exposure to diverse viewpoints. Furthermore, these systems tend to reinforce recognition disparities between prominent and lesser-known researchers and may inadvertently contribute to the dissemination of misinformation  (Polonioli, 2021; Bolaños et al., 2024). To address these biases, recommendation algorithms can be enhanced by emphasizing content-based rather than author-based recommendations and by incorporating robust evaluation mechanisms to strengthen the credibility of suggested materials.
In contrast, AI-driven hypothesis generation presents more pronounced ethical challenges. First, the attribution of intellectual property rights and authorship for AI-generated hypotheses remains ambiguous  (Majumder et al., 2024a). Additionally, the widespread generation of low-quality content poses a risk of diluting the integrity of the academic landscape (Hu et al., 2024a), while the potential misuse of such technologies for illicit purposes cannot be overlooked  (Si et al., 2024). Addressing these concerns necessitates the development of robust accountability frameworks, the assignment of clear responsibility for AI-generated outputs to researchers, and the establishment of appropriate legal and regulatory mechanisms.
During the hypothesis validation phase, automated systems for scientific fact-checking remain underdeveloped. This limitation may be exploited by malicious actors to create advanced misinformation generators capable of circumventing existing fact-checking tools  (Wadden et al., 2022b). Likewise, in the context of experimental validation, there is a risk of unethical or legally questionable experiments being designed  (Eger et al., 2025). These concerns underscore the need for continued research into model safety.
During the manuscript publication stage, several challenges remain. Text generated by AI models may carry a risk of plagiarism  (Salvagno et al., 2023; Gupta and Pruthi, 2025), while AI-assisted peer reviews often offer vague feedback and exhibit inherent biases  (Schintler et al., 2023; Drori and Te’eni, 2024; Pataranutaporn et al., 2025). To address these issues, the development of robust detection methods is essential. However, current detection tools are still in the early stages of maturity  (Gupta and Pruthi, 2025).
Appendix D Ability Comparison
An effective survey should not only summarize existing methods within a field but also provide comparative analyses of different approaches. However, the domain of AI for Research remains in its early stages, with many areas lacking standardized benchmarks and even established evaluation metrics. To facilitate a clearer understanding of the distinctions among various methods, we draw on existing literature (Kang et al., 2023; Bolaños et al., 2024; Luo et al., 2025; Vladika and Matthes, 2023; Yang et al., 2023c; Li and Ouyang, 2022, 2024; Lin et al., 2023a) and adopt attribute graphs to compare representative approaches within each subfield, as illustrated in table  §1 to table  §8.
Method Human-Computer Interaction LLM Required Information Return Information Relevance Source
ComLittee (Kang et al., 2023) - Authorship Graphs Meta data with relevant authors R, Co, Ci
ArZiGo (Pinedo et al., 2024) - User Interest Meta data R
PaperWeaver (Lee et al., 2024) Collected Papers Meta data with description R
Kang et al. (2022) - - Author’s social network relationships +Reference relationship Meta data with relevant authors R
Table 1: Research Paper Recommendation, we referred to  Kang et al. (2023) for comparing different methods, where R represents Paper recommender score, Co represents Co-author relationship, and Ci represents Cited author relationship.
Method Research Field Across Stages Human Interaction Task Input Output Evaluation Method
AutoSurvey (Wang et al., 2024e) Any - Outline Generation, +Full-text Generation Title & Full Content Literature Survey LLM & Human
CHIME (Hsu et al., 2024) Biomedicine - Outline Generation Title & Full Content Hierarchical Outline Automatic Metrics
Knowledge Navigator (Katz et al., 2024) Any - - Outline Generation Title & Full Content Hierarchical Outline LLM & Human
Relatedly (Palani et al., 2023) Any - - Full-text Generation Title & Related Work Literature Survey Human
STORM (Shao et al., 2024) Any - - Outline Generation, +Full-text Generation Title & Full Content Literature Survey LLM & Automatic Metrics
Table 2: Scientific Literature Review, we referred to Bolaños et al. (2024) and made modifications, thereby comparing different methods.
Method Research Field Across Stages Human Interaction Multi-agent Trained Model Online RAG Novelty Feasibility Validity
COI (Li et al., 2024a) Any - - -
Learn2Gen (Li et al., 2024c) Artification Intelligence - - -
MatPilot (Ni et al., 2024) Materials Science - - -
SciAgents (Ghafarollahi and Buehler, 2024) Any - - - -
SciMON (Wang et al., 2024c) Any - - - - - -
Table 3: Hypothesis Generation, we referred to Luo et al. (2025) and made modifications, thereby comparing different methods.
Method Input Document Retrieval Human Interaction Rationale Selection Evidence Format Output
MULTIVERS (Wadden et al., 2022b) Claim & scientific abstract Provided - Longformer Document Label & sentence-level rationales
SFAVEL (Bazaga et al., 2024) Claim Pre-trained Language Model - - knowledge graph Top-K Facts & Corresponding Relevance Scores
ProToCo (Zeng and Gao, 2023) Claim-Evidence Pair Provided - - Sentence Label
MAGIC (Kao and Yen, 2024) Claim Provided - Dense Passage Retriever Sentence Label
aedFaCT (Altuncu et al., 2023) News Article Google Search Human Document Evidence
Table 4: Scientific Claim Verification, we referred to Vladika and Matthes (2023) and made modifications, thereby comparing different methods.
Method Generation Based Stepwise Heuristic Search Informal or Formal Human-authored Realistic Proof
IBR (Qu et al., 2022) - Informal -
GPT-f (Polu and Sutskever, 2020) - Formal
DT-Solver (Wang et al., 2023b) Formal
POETRY (Wang et al., 2024a) - - Formal
Table 5: Theorem proving, we referred to Yang et al. (2023c) and made modifications, thereby comparing different methods.
Method Research Field Across Stages Human Interaction Multi-agent Task Input External tools
AutoML-GPT (Zhang et al., 2023) Artification Intelligence - - - Automated Machine Learning Task-oriented Prompts -
Chemcrow (Bran et al., 2024) Chemistry - - Chemical Task Task Description
DOLPHIN (Yuan et al., 2025) Any - Automated Scientific Research Idea
MechAgents (Ni and Buehler, 2023) Physics - - Mechanical Problem - -
Manning et al. (2024) Social Science - Simulating Human - -
Table 6: Experiment Validation: we use attribute diagrams to compare different schemes, and the table design refers to Hypothesis Generation.
Method Across Stages Human Interaction Task Input Evaluation Method
AI Scientist (Lu et al., 2024) - Full-text Generation Manuscript Template & Experimental Results & Hypothesis LLM
data-to-paper (Ifargan et al., 2024) Full-text Generation Experimental Results & Hypothesis -
ScholaCite (Martin-Boyle et al., 2024) - - Related Work Generation Title & Abstract & Citation Citation Graph Metrics
SciLit (Gu and Hahnloser, 2023) - Citation Generation Keywords Automatic Metrics
Gu and Hahnloser (2024) - - Citation Generation Citation Intent & Keywords Human
Table 7: Manuscript Writing, we referred to Li and Ouyang (2022, 2024) and made modifications, thereby comparing different methods.
Method Across Stages Human Interaction Paper Review Meta Review Multi-agent Trained Model Output
Gamma-Trans (Muangkammuen et al., 2022) - - - - Peer-review Score
MARG (D’Arcy et al., 2024a) - - - - Peer-review Comments
CycleResearcher (Weng et al., 2024) - - - Peer-review Comments & Score
PeerArg (Sukpanichnant et al., 2024) - - - - - Final Decision
GLIMPSE (Darrin et al., 2024) - - - - Summary of Peer-review
Table 8: Peer Review, we referred to Lin et al. (2023a) and made modifications, thereby comparing different methods.
Task Benchmark Domain Size Input Output Metric
SCHOLAT Li et al. (2020) Research Paper Recommendation 34,518 - - Recall, Precission, F1-score
ACL selection network Tao et al. (2020) Research Paper Recommendation 18,718 Topics Related Papers Accuracy
CiteSeer Kang et al. (2021) Research Paper Recommendation 1,100 Paper Related Papers Correlation Coefficient
SciReviewGen Kasanishi et al. (2023) Systematic Literature Review 10,000+ Abstracts literature review ROUGE
FacetSum Meng et al. (2021) Systematic Literature Review 60,024 Source Text+Facet Summary of Facet ROUGE
BigSurvey Liu et al. (2022) Systematic Literature Review 7,000+ Abstracts Survey Paragraph ROUGE, F1-score
SCHOLARQABENCH Asai et al. (2024) Systematic Literature Review 2,200 Question Answer with Citations Accuracy, Coverage, Citations + Relevance, Usefulness
HiCaD Zhu et al. (2023) Systematic Literature Review 7,600 Reference Papers Catalogues Catalogue Edit Distance Similarity (CEDS) + Catalogue Quality Estimate (CQE)
CLUSTREC-COVID Katz et al. (2024) Systematic Literature Review 2,284 Titles, Abstracts Topic Clusters per Topic
CHIME Hsu et al. (2024) Systematic Literature Review 2,174 Topic Hierarchies F1-score
Tian et al. (2024) Systematic Literature Review 700 Subject, Reference Title,Content -
MASSW Zhang et al. (2024c) Hypothesis Generation 152000 Context of Literature Hypothesis BLEU, ROUGE, BERTScore, + Cosine Similarity, BLEURT
IdeaBench Guo et al. (2024) Hypothesis Generation 2,374 Instruction, Background Information Hypothesis Insight Score, BERTScore, Novelty, + LLM Similarity Rating, Feasibility
SCIMON Wang et al. (2024c) Hypothesis Generation - Background Context Idea ROUGE, BERTScore +BARTScore, Novelty
MOOSEYang et al. (2024b) Hypothesis Generation 50 Background, Inspiration Hypothesis Validness, Novelty + Helpfulness
DISCOVERYBENCH Majumder et al. (2024b) Hypothesis Generation 1,167 Data Discovery Hypothesis Match Score
Hypothesis LiveIdeaBench Ruan et al. (2024a) Hypothesis Generation - Scientific Keywords Idea Originality, Feasibility + Fluency, Flexibilit
Formulation Kumar et al. (2024) Hypothesis Generation 100 Paper without Future Work Idea Idea Alignment Score, Idea Distinctness Index
SciRIFF Wadden et al. (2024) Scientific Claim Verification 137,000 Evidence, Task prompt Structured Paragraph F1, BLEU
SCIFACT Wadden et al. (2020) Scientific Claim Verification 1,409 Claim, Evidence Rationale Sentences, Label Precision, Recall, Micro-F1
SCIFACT-OPEN Wadden et al. (2022a) Scientific Claim Verification 279 Claim, Evidence Rationale Sentences, Label Precision, Recall,Micro-F1
MISSCI Glockner et al. (2024b) Scientific Claim Verification 435 Claim, Premise, Context Verification Micro F1-score,P@1,Arg@1 + METEOR Score,BERTScore +NLI-A, NLI-S, Matches@1
FEVER Thorne et al. (2018) Scientific Claim Verification 185,445 Claim, Evidence Label, Necessary Evidence F1-Score,Oracle Accuracy + Accuracy,Recall
XClaimCheck Kao and Yen (2024) Scientific Claim Verification 16,177 Claim, Evidence Label, Argument Macro-F1, Accuracy
HEALTHVER Sarrouti et al. (2021) Scientific Claim Verification 14330 Claim, Evidence Label Macro Precision, Macro Recall + Macro F1-score, Accuracy
QuanTemp V et al. (2024) Scientific Claim Verification 15,514 Claim, Evidence Label Weighted-F1 Score, Macro-F1, BLEU, + BERTScore, Cohen’s Kappa Score + Human Evaluation
SCITAB Lu et al. (2023) Scientific Claim Verification 1,225 Claim, Evidence Label Macro-F1
Check-COVID Wang et al. (2023a) Scientific Claim Verification 1,504 Claim Evidence Accuracy, Precision, Recall, Macro-F1
HealthFC Vladika et al. (2024) Scientific Claim Verification 750 Claim, Evidence Label Precision, Recall, F1-Macro
FACTKG Kim et al. (2023) Scientific Claim Verification 108,000 Claim, Evidence Label Accuracy
BEAR-FACT Wührl et al. (2024a) Scientific Claim Verification 1,448 Claim, Evidence +Entity/Relation Information Label F1-Score
MINIF2F Zheng et al. (2022) Theorem Proving 488 Problem, Theorem Proof Pass Rate
FIMO Liu et al. (2023a) Theorem Proving 149 Problem, Theorem, statements Proof Pass Rate
LeanDojo Yang et al. (2023a) Theorem Proving 98,734 Problem, Theorem Proof R@k, MRR, Pass Rate
Lean-github Wu et al. (2024) Theorem Proving 28,597 Problem, Theorem Proof Accuracy, Pass Rate
TRIGO-real Xiong et al. (2023) Theorem Proving 427 Problem, Theorem Proof Pass Rate, Accuracy, EM@n
TRIGO-web Xiong et al. (2023) Theorem Proving 453 Problem, Theorem Proof Pass Rate, Accuracy, EM@n
TRIGO-gen Xiong et al. (2023) Theorem Proving - Problem, Theorem Proof Pass Rate, Accuracy, EM@n
CoqGym Yang and Deng (2019) Theorem Proving 71,000 Problem, Theorem Proof Success Rate
MLAgentBench Huang et al. (2024b) Experiment Validation 13 - - Competence, Efficiency
AAAR-1.0 Lou et al. (2024) Experiment Validation - Instance, Papers Design, Explanation S-F1, S-Precision, S-Recall + S-Match, ROUGE
TASKBENCH Shen et al. (2024) Experiment Validation 17,331 - - ROUGE, t-F1, v-F1 +Normalized Edit Distance
Spider2-V Cao et al. (2024a) Experiment Validation 494 Task Experiment Execution Success Rate
CORE-Bench Siegel et al. (2024) Experiment Validation 270 Task Requirements Experiment Result Accuracy
LAB-Bench Laurent et al. (2024) Experiment Validation 2400 Multiple-choice Question Answer Accuracy, Precision, Coverage
PaperBench Starace et al. (2025) Experiment Validation 20 Paper, Additional Information Code Replication Score
Hypothesis SUPER Bogin et al. (2024) Experiment Validation 801 Task Requirements - Accuracy, Landmark-Based Evaluation
Validation ScienceAgentBench Chen et al. (2025) Experiment Validation 102 Task Instruction, Dataset Information +Expert-Provided Knowledge Program Valid Execution Rate, Success Rate, CodeBERTScore, API Cost
SciCap+ Yang et al. (2023b) Manuscript Writing 414,000 Figure, OCR tokens + Mention Paragraph Caption BLEU, ROUGE, METEOR + CIDEr, SPICE
AAN Corpus Radev et al. (2013) Manuscript Writing - - - -
SciSummNet Yasunaga et al. (2019) Manuscript Writing 1,000 Paper,Citation Sentence Summary ROUGE
CiteBench Funkquist et al. (2023) Manuscript Writing 358,765 Cited Papers, Context Citation Text ROUGE, BERTScore
ALCE Gao et al. (2023b) Manuscript Writing 3,000 Question Answer with Citations Recall, Precision
GCite Wang et al. (2022b) Manuscript Writing 2,500 Citing/Cited Paper Citation Text BLEU, ROUGE
ARXIVEDITS Jiang et al. (2022b) Manuscript Writing 1,000 Sentence Pairs Sentence, Intent Precision,Recall,F1-score
CASIMIR Jourdan et al. (2024) Manuscript Writing 15,646 Original Sentence Revised Sentence Exact-match (EM),SARI, BLEU, + ROUGE-L,Bertscore
ParaRev Jourdan et al. (2025) Manuscript Writing 48,203 Original Paragraph Revised Paragraph ROUGE-L,SARI + BertScore
SCHOLAWRITE Wang et al. (2025) Manuscript Writing 62,000 Before-text Writing Intention, After-text F1-score, Lexical Diversity, Topic Consistency, Intention Coverage
MReD Shen et al. (2022) Peer Review 7,089 Reviews Meta-Review ROUGE
ORSUMZeng et al. (2024) Peer Review 15,062 Reviews Meta-Review ROUGE-L, BERTScore, FACTCC + SummaC, DiscoScore
PeerRead v1 Kang et al. (2018) Peer Review 107,000 Reviews Accept/Reject Accuracy
NLPeer Dycke et al. (2023) Peer Review 5,000 Reviews,Paper Review Score, Connection, + Review Category MRSE, F1-macro + Precision, Recall
AMPERE Hua et al. (2019) Peer Review 400 Review Review with Type Precision, Recall, F1-score
MOPRD Lin et al. (2023b) Peer Review 6,578 Reviews,Paper Editorial Decision, Review, + Meta-Review, Author Rebuttal ROUGE, BARTScore
ARIES D’Arcy et al. (2024b) Peer Review 1,720 Review Comment, Edits Comment-Edit Pairs Precision, Recall, F1-score
ASAP-Review Yuan et al. (2022) Peer Review - Paper Review Aspect Coverage, Aspect Recall, +Semantic Equivalence +Human: Recommendation Accuracy(RAcc), +Informativeness(Info),Aspect-level, +Constructiveness(ACon) and Summary accuracy
ReviewMT Tan et al. (2024) Peer Review 26,841 Paper Review Dialogue ROUGE,BLEU,METEOR
Manuscript ReAct Choudhary et al. (2021) Peer Review 6,250 Review Classification of Review Accuracy
Publication PEERSUM Li et al. (2023) Peer Review - Reviews Meta-Review ROUGE,BERTScore,UniEval,ACC
Table 9: An overview of benchmarks on AI for research. In the Input, Output, and Metric columns, the ’+’ symbol indicates that the row is a continuation of the previous row.
Tool Research Paper Recommendation Systematic Literature Review Hypothesis Generation Scientific Claim Verification Theorem Proving Experiment Verification Manuscript Writing Peer Review Reading Assistance
Connected Paper $\checkmark$
Inciteful $\checkmark$
Litmaps $\checkmark$
Pasa $\checkmark$
Research Rabbit $\checkmark$
Semantic Scholar $\checkmark$ $\checkmark$
GenGO $\checkmark$ $\checkmark$
Jenni AI $\checkmark$ $\checkmark$ $\checkmark$
Elicit $\checkmark$ $\checkmark$
Undermind $\checkmark$ $\checkmark$
OpenScholar $\checkmark$ $\checkmark$
ResearchBuddies $\checkmark$ $\checkmark$
Hyperwrite $\checkmark$ $\checkmark$ $\checkmark$
Concensus $\checkmark$ $\checkmark$ $\checkmark$
Iris.ai $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MirrorThink $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SciSpace $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AskYourPDF $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Iflytek $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
FutureHouse $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Enago Read $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Aminer $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OpenRsearcher $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ResearchFlow $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
You.com $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GPT Researcher $\checkmark$
PICO Portal $\checkmark$
SurveyX $\checkmark$
Scinence42:Dora $\checkmark$ $\checkmark$
STORM $\checkmark$ $\checkmark$
ChatDOC $\checkmark$ $\checkmark$
Scite $\checkmark$ $\checkmark$
Silatus $\checkmark$ $\checkmark$
Agent Laboratory $\checkmark$ $\checkmark$ $\checkmark$
Sider $\checkmark$ $\checkmark$ $\checkmark$
Quillbot $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Scholar AI $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AI-Researcher $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AI Scientist $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Isabelle $\checkmark$
LeanCopilot $\checkmark$
Llmstep $\checkmark$
Proverbot9001 $\checkmark$
chatgpt_academic $\checkmark$
gpt_academic $\checkmark$
HeadlineAnalyzer $\checkmark$
Langsmith Editor $\checkmark$
Textero.ai $\checkmark$
Wordvice.AI $\checkmark$
Writesonic $\checkmark$
Writefull $\checkmark$ $\checkmark$
Covidence $\checkmark$
Penelope.ai $\checkmark$
Byte-science $\checkmark$
Cool Papers $\checkmark$
Explainpaper $\checkmark$
Uni-finder $\checkmark$
Table 10: Tools for Research Paper Assistance
{forest}
forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, minimum height=2.5em, minimum width=5em, edge+=darkgray, line width=1pt, s sep=6pt, inner xsep=0.6em, inner ysep=0.9em, line width=0.8pt, text width=10em, ver/.style= rotate=90, child anchor=north, parent anchor=south, anchor=center, text width=13em , leaf/.style= text opacity=1, fill opacity=.5, fill=hidden-blue!90, text=black, text width=50em, font=, inner xsep=0.6em, inner ysep=0.9em, draw, , , [ A survey of AI for Research, ver [ Hypothesis
Formulation  (§2) [ Knowledge Synthesize  (§2.1) [ Research Paper
Recommendation  (§2.1.1) [ PaperWeaver Lee et al. (2024), ArZiGo Pinedo et al. (2024), ComLittee Kang et al. (2023),  Stergiopoulos et al. (2024),  Kang et al. (2022),  Kreutz and Schenkel (2022) ,  Shahid et al. (2020),  Bai et al. (2019),  Li and Zou (2019),  Beel et al. (2016) , leaf, text width=50em ] ] [ Systematic
Literature Review  (§2.1.2) [ AutoSurvey  Wang et al. (2024e) , OpenScholar  Asai et al. (2024) , HiReview  Hu et al. (2024b), PaperQA2  Skarlinski et al. (2024), Knowledge Navigator  Katz et al. (2024), CHIME  Hsu et al. (2024), STORM  Shao et al. (2024), LitLLM  Agarwal et al. (2024a), Bio-SIEVE  Robinson et al. (2023), KGSum  Wang et al. (2022a), Agarwal et al. (2024b), Lai et al. (2024), Susnjak et al. (2024), Bolaños et al. (2024), Block and Kuckertz (2024), Zhu et al. (2023) , Altmami and Menai (2022) , leaf, text width=50em ] ] ] [ Hypotheses Generation  (§2.2) [ Other Works [ MOOSE-Chem  Yang et al. (2024c), IdeaSynth  Pu et al. (2024), Si et al. (2024), Kumar et al. (2024), Girotra et al. (2023), Qi et al. (2023), Krenn et al. (2022), Henry and McInnes (2017) , leaf, text width=50em ] ] [ Input Data Quality [ SciPIP Wang et al. (2024d), COI Li et al. (2024a), SciAgents Ghafarollahi and Buehler (2024), DATAVOYAGER  Majumder et al. (2024a), SciMON Wang et al. (2024c), Buehler (2024), Liu et al. (2024a) , leaf, text width=50em ] ] [ Hypothesis Quality [ Agentrxiv Schmidgall and Moor (2025) Dolphin Yuan et al. (2025), VIRSCI Su et al. (2024), Nova Hu et al. (2024a), AI Scientist Lu et al. (2024), SGA Ma et al. (2024), HypoGeniC  Zhou et al. (2024), ResearchAgent Baek et al. (2024), Acceleron Nigam et al. (2024), MOOSE  Yang et al. (2024b) , leaf, text width=50em ] ] ] ] [ Hypothesis Validation  (§3) [ Scientific Claim Verification  (§3.1) [ Claim [ HiSS Zhang and Gao (2023), SFAVEL Bazaga et al. (2024), ProToCo Zeng and Gao (2023), Wührl et al. (2024a) , leaf, text width=50em ] ] [ Evidence [ Glockner et al. (2024a), Vladika and Matthes (2024b), Vladika and Matthes (2024a), Wührl et al. (2024b), Pan et al. (2023b), Glockner et al. (2022), Wadden et al. (2022b) , leaf, text width=50em ] ] [ Verification [ ClaimVer Dammu et al. (2024), MAGIC Kao and Yen (2024), FactKG Kim et al. (2023), aedFaCT Altuncu et al. (2023), PROGRAMFC Pan et al. (2023a), Zhang et al. (2024b), Muharram and Purwarianti (2024), Eldifrawi et al. (2024), Cao et al. (2024b), Das et al. (2023), Wu et al. (2023) , leaf, text width=50em ] ] ] [ Theorem Proving  (§3.2) [ DeepSeek-Prover Xin et al. (2024), Lean-STaR Lin et al. (2024), Thor Jiang et al. (2022a), COPRA Thakur et al. (2024), Logo-power Wang et al. (2024b), Baldur First et al. (2023), Mustard Huang et al. (2024c), DT-Solver Wang et al. (2023b), HTPS Lample et al. (2022), GPT-f Polu and Sutskever (2020), Wang et al. (2024a), Song et al. (2024), Li et al. (2024e), Zhao et al. (2023b), Yang et al. (2023c), Jiang et al. (2023) , leaf, text width=62.5em ] ] [ Experiment Validation  (§3.3) [ MatPilot Ni et al. (2024), CRISPR-GPT Huang et al. (2024a), AgentHPO  Liu et al. (2024c), AutoML-GPT Zhang et al. (2023), MLCopilot Zhang et al. (2024a), Arlt et al. (2024), Kambhampati et al. (2024), Ruan et al. (2024b), Huang et al. (2024b), Boiko et al. (2023), Bran et al. (2024), Szymanski et al. (2023), Liu et al. (2023b), Manning et al. (2024), Mou et al. (2024) , leaf, text width=62.5em ] ] ] [ Manuscript
Publication  (§4) [ Manuscript Writing  (§4.1) [ Citation Text
Generation [ SciLit Gu and Hahnloser (2023), DisenCite Wang et al. (2022b), Gu and Hahnloser (2024), Gao et al. (2023b), Yu et al. (2022), Jung et al. (2022) , leaf, text width=50em ] ] [ Related Work
Generation [ ScholaCite Martin-Boyle et al. (2024), UR3WG Shi et al. (2023), Li and Ouyang (2024), Yu et al. (2024b), Nishimura et al. (2024), Li and Ouyang (2022) , leaf, text width=50em ] ] [ Complete Manuscripts
Generation [ Cocoa Feng et al. (2024), Step-Back Tang et al. (2024a), data-to-paper Ifargan et al. (2024), ARIES D’Arcy et al. (2024b), OREO Li et al. (2022), R3 Du et al. (2022a), Lai et al. (2024), Chamoun et al. (2024), Pividori and Greene (2024), Jourdan et al. (2023), Lin (2024) , leaf, text width=50em ] ] ] [ Peer Review  (§4.2) [ Paper Review
Generation [ CycleResearcher Weng et al. (2024), OpenReviewer Idahl and Ahmadi (2024) RelevAI-Reviewer Couto et al. (2024), AgentReview Jin et al. (2024), SWIF2T Chamoun et al. (2024), MARG D’Arcy et al. (2024a), Quality Assist Setio and Tsuchiya (2022), KID-Review Yuan and Liu (2022) Tan et al. (2024), Liang et al. (2023), Robertson (2023), Muangkammuen et al. (2022), , leaf, text width=50em ] ] [ Meta-Review
Generation [ Deeprview Zhu et al. (2025) PeerArg Sukpanichnant et al. (2024) GLIMPSE Darrin et al. (2024), MetaWriter Sun et al. (2024), CGI2 Zeng et al. (2023), Li et al. (2024b), Santu et al. (2024), Kumar et al. (2023), Li et al. (2023) , leaf, text width=50em ] ] ] ] ]
Figure 6: Taxonomy of Hypothesis Formulation, Hypothesis Validation and Manuscript Publication (Full Edition).